Prove that if x & y are odd positive integers then x^2+y^2 is even but not divisible by 4

Grade:10

1 Answers

S. Haris Ahmed Irfan
9 years ago
all odd positive integers are of form 2q+1 therefore let x=2q+1 and y=2p+1 (where p and q are integers) x^2=(2q+1)^2=4q^2+1+4q y^2=(2p+1)^2=4p^2+1+4p therefore x^2+y^2=4q^2+1+4q+4q^2+1+4q after rearranging 4(q^2+p^2+p+q)+2 this number is clearly even but is not divisible by 4 {since 2 is added at the end}